OpenStudy (anonymous):

Calculate the future value when P=$7,500, R=9%, compounded monthly, and t=8 years. Round to the nearest cent.

OpenStudy (anonymous):

I got r=.09 and P*R*T but not getting it still?

OpenStudy (tkhunny):

COMPOUND INTEREST R = 0.09 -- The annual accumulation interest rate to be compounded monthly. i = R/12 = 0.0075 -- The monthly interest rate. s = 1+i = 1.0075 -- Monthly Accumulation Factor \(7500\cdot s^{8*12}\)

OpenStudy (anonymous):

OH boy you still there?

OpenStudy (tkhunny):

Did you do it?

OpenStudy (anonymous):

so its 7500 and times 8 to 96 power????

OpenStudy (tkhunny):

8? No.. That's an "s". See the definition just above.

OpenStudy (anonymous):

OH sorry thought it was a 8...so no havent got it yet

OpenStudy (anonymous):

7,961.99????

OpenStudy (anonymous):

is that right??

OpenStudy (tkhunny):

What value did you use for "s"?

OpenStudy (anonymous):

1.0075

OpenStudy (anonymous):

or is it 1 + 0.0075

OpenStudy (anonymous):

Oh gosh Im confusing myself lol

OpenStudy (tkhunny):

I think not. You entered the interest rate and divided it by 12 a second time. That's no good. Whatever you did, use the 9% and it should come out.

OpenStudy (anonymous):

Oh dear gosh I wish I was good at Math this is crazy.... so its 7500 * 1.0075* 0.09???? Im sorry I suck at math ...just this problem and Im done with this worksheet

OpenStudy (tkhunny):

Stop! Take a deep breath. Slow down. How are you trying to do this. Calculator? Spreadsheet? Financial calculator? Paper and pencil? Just what is it you are doing?

OpenStudy (anonymous):

LOL ok sorry been doing this since 3 this afternoon Im burning out.... Im using both paper and calculator on computer thats it

OpenStudy (tkhunny):

Does it have "scientific" functions? Most interesting would be one of these: \(x^{y}\) -- an exponent button.

OpenStudy (anonymous):

No I wish It would help me out

OpenStudy (anonymous):

Thats why its taking me so long I have to do it all on paper

OpenStudy (tkhunny):

We can use the calculator if it has a memory. How about \(M+\;or\;MS\) or something like that.

OpenStudy (anonymous):

Whatever smartscore means up here Im not feelin it LOL.... it has M+ M- stuff like that

OpenStudy (tkhunny):

Okay, let's run a little test and then we'll be done in a jiffy. Press these keys (or click buttons) 2 x = Does it give 4 as the answer?

OpenStudy (anonymous):

Ok let me try

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

One more test. 2 x = x = Just hit the "2" once and the the sequence "x =" twice. Does it give 16 as the result?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Every time you hit the "x =" sequence, you SQUARED the number in the display. We are about to use this feature to our advantage. Listen very carefully. Type in the accumulation factor, 1.0075 Hit the "x = " sequence FIVE TIMES That's 1.0075 x = x = x = x = x = and tell save the result or write it down here so I can check it out.

OpenStudy (anonymous):

1.27011124

OpenStudy (tkhunny):

Perfect. Now hit "x =" one more time - with that number on the display. Tell me the result.

OpenStudy (anonymous):

1.61318256

OpenStudy (tkhunny):

Exactly correct. One more thing, and then I'll explain what we just did. Multiply that number by 1.27011124 and tell me the result.

OpenStudy (anonymous):

it wont let me put the last 4 in?????

OpenStudy (tkhunny):

That's fine. Let's just see how close we can get.

OpenStudy (anonymous):

2.04892116

OpenStudy (tkhunny):

That's it. Exactly correct (well, with the silly 4 missing, anyway.) Explanation: When we hit "x =" five times, we calculated \(1.0075^{32}\). When we hit "x =" one last time, squaring the previous result, we created \(1.0075^{64}\). When we multiplied them together, we managed \( 1.0075^{32}\cdot 1.0075^{64} = 1.0075^{96}.\) Exactly what we needed. Multiply that by 7500 and we'll be done!

OpenStudy (anonymous):

15,366!!!!!!! :) That is awesome!!!!

OpenStudy (anonymous):

wow you made is so simple ha. :)

OpenStudy (tkhunny):

-) Now, go get a better calculator. That was way harder than it should be! Your phone probably has a scientific calculator app.

OpenStudy (anonymous):

Yes I agree I do need one really bad!!

OpenStudy (anonymous):

I do have more problems that no one on here can answer I mean no one I gave up...I dont want to bother you anymore but if you have time tomorrow or if you have time now up to you...they are stumping everyone and you seem amazing with math! :)

OpenStudy (anonymous):

I have got to get me a new calculator!!!!! :)

OpenStudy (anonymous):

Its on another worksheet :)

OpenStudy (anonymous):

I wanted to also thank you for your time! :) means alot!

OpenStudy (tkhunny):

They can be downloaded http://translateye.com/download/calculator.php or just used online. No worries. It's what we do around here!

OpenStudy (anonymous):

OH thank you!!!! Im downloading it!!!!

OpenStudy (tkhunny):

We can get you through it. This is where I live some of the time. :-)

OpenStudy (anonymous):

LOL oh good Im so glad :) did you want to help me tonight or tomorrow I dont know where you live maybe to late for you :)

OpenStudy (tkhunny):

East Coast USA - Past my bedtime! ZZZZZzzzz......

OpenStudy (anonymous):

LOL I understand thats why I was checking...I will look for you on here tomorrow night then :) I really do appreciate all of your time :) I should go to bed too almost 11 :)

OpenStudy (anonymous):

Thanks again!!!! Have a wonderful night!!!!!! :0)

OpenStudy (tkhunny):

There's the online one I was looking for! http://web2.0calc.com/

OpenStudy (anonymous):

OH THANK YOU!!!! have it saved to my desktop!!!!! Thanks alot!!!!!

OpenStudy (anonymous):

Nighty Night :)